Find all x belongs to R that satisfy the following inequalities: i)|x|+|x+1|<2 ii)|x^2-1|<_3


1
Expert's answer
2021-12-13T16:28:05-0500

i)

a)

"x<-1:"


"-x-x-1<2"

"x>-\dfrac{3}{2}"

"x\in(-\dfrac{3}{2}, -1)"


b)

"-1\leq x<0:"


"-x+x+1<2"

"2>1"

"x\in[-1, 0)"


c)

"x\geq 0"


"x+x+1<2"

"x<\dfrac{1}{2}"

"x\in[0, \dfrac{1}{2})"

Answer: "(-\dfrac{3}{2}, \dfrac{1}{2})"


ii)


"|x^2-1|\leq3"

"-3\leq x^2-1\leq 3"

"-2\leq x^2\leq4"

"-2\leq x\leq 2"

Answer: "[-2, 2]"